# RhythmFormer: Extracting rPPG Signals Based on Hierarchical Temporal Periodic Transformer

### v1

@@ 3.3 Hierarchical Temporal Periodic Transformer
Past studies [14, 16, 7] have demonstrated the effectiveness of modeling periodic tasks with multiple temporal scales, primarily extracting temporal information through concatenation from paths at various scales. In the context of rPPG tasks, similar to many spatial multi-scale models [25, 43], we believe that a pyramid-shaped hierarchical multi-scale learning can facilitate the provision of information from earlier learned scales to later learned scales. The proposed Hierarchical Temporal Periodic Transformer consists of three stages. It first reinforces inter-frame rPPG features at a smaller time scale and then models periodic features at a larger time scale, effectively reducing noise interference. Each stage consists of a TPT Block with a sampling coefficient of $n$, adjusting the specific time scale through n times sampling modules within the block.
The TPT block first aggregates highly correlated regions through the pre-attention stage, and then extracts features with the refined attention stage at a finer granularity. Specifically, for the TPT block input $X_{in}\in\mathbb{R}^{C\times T\times H/16\times W/16}$, we first perform $n$ rounds of temporal downsampling, where the temporal information is aggregated at a granularity of $2^{n}$ without overlap. The downsampling module consists of a BN layer and a 3D convolution layer with the kernel and stride size of $(2\times 1\times 1)$, resulting in an output $X_{sample}\in\mathbb{R}^{C\times{T}/{2^{n}}\times H/16\times W/16}$.
@@
The temporal difference convolution (TDC) [55] has been proven to effectively capture fine-grained local temporal difference features [53, 60, 56, 57]. We utilize the TDC module for projecting Q and K, while V is projected using linear projection. The projected Q and K are subjected to a 3D average pooling operation to compute attention within a larger spatio-temporal receptive field, which filters out low-correlation key-value pairs. Specifically, for input $Q,K\in\mathbb{R}^{C\times{T}/{2^{n}}\times H/16\times W/16}$, the pooling convolutional kernel size is $(T^{\prime}\times H^{\prime}\times W^{\prime})$.
$T^{\prime}=[\frac{T}{T_{s}}],H^{\prime}=[\frac{H/16}{4}],W^{\prime}=[\frac{W/16}{4}]$ (4)
@@
Where $T_{s}=MAX(2^{x},2^{n})$, and $x$ is the partition coefficient. In the experiments, since the frame rate of the dataset is uniformly downsampled to 30, $x$ is set to 2. The number of tokens in the pre-attention is $T_{s}\times 4\times 4$. The scores of the pre-attention are computed as follows:
$Score\ =\ Q{{}^{\prime}(K^{\prime})}^{T}$ (5)
@@
The pre-attention stage is designed to obtain the indices of top-k positions with high correlation, to filter out low-correlation key-value pairs. The pre-attention captures spatio-temporal correlations within a large receptive field, gathering the regions highly correlated with Q in both K and V. If given a token of the skin area as a query, it corresponds to gathering skin regions with similar temporal phases to the Query. The gathered key and value are expressed as:
$K^{g}\ =\ gather(K,topkIndex(Score))$ (6)
$V^{g}\ =\ gather(V,topkIndex(Score))$
@@
The refined attention stage is computed on the top-k (high-correlation) regions identified after the pre-attention gathering. Subsequently, the local positional information is enhanced through a LCE module following self-attention [35, 61]. The $i_{th}$ self-attention head output ${SA}_{i}$ is:
${SA}_{i}\ =Softmax(\ Q{(K^{g})}^{T}/\sqrt{D})\ V^{g}\ +\ LCE(V)$ (7)
@@
where $D$ represents the dimensionality of single-head attention, and LCE is a 3D convolution with a kernel size of $(3\times 3\times 3)$. The output of multi-head self-attention is obtained by concatenating the outputs from all attention heads, followed by a linear projection through the weight matrix $W\in\mathbb{R}^{D\times D}$.
$MHSA=Concat(SA_{1},SA_{2},...,SA_{h})W$ (8)
@@
Like vanilla transformer [44], residual connection, BN, and feed-forward would be conducted after MHSA. Diverging from that, inspired by [56, 57], an additional hidden layer is incorporated into the feed-forward layer. This hidden layer consists of a three-dimensional convolution, BN, and GELU, intended to refine local inconsistency and reinforce relative position cues. Finally, the Temporal Periodic Transformer will generate output through $n$ temporal upsampling modules. The dimensionality of the output $X_{out}$ will be restored to match the consistency of the input $X_{in}$, $X_{out}\in\mathbb{R}^{C\times T\times H/16\times W/16}$.
3.4 HR Hybrid Loss
Previous loss functions have constrained rPPG signals from both the temporal domain [3, 51, 23, 37] and the frequency domain [32, 53, 39]. Temporal domain constraints primarily operate at the signal-trend level, e.g. the amplitude of the signal. Frequency domain constraints, on the other hand, focus on guiding the learning of periodic features. However, due to the potential mixture of information and noise in the ground truth BVP, there is a gap between the BVP loss and HR metrics during the training process. Therefore, we introduce a loss function that characterizes the HR distance, which is a non-gradient-bearing loss used solely to guide the network in selecting the optimal checkpoint.
Since the collection of ground truth BVP may involve errors, instead of treating HR as a deterministic value, HR is viewed as a probability distribution centered around the HR value. The HR distance is characterized by the distance between the ground truth HR distribution and the predicted HR distribution. Specifically, the HR distribution is expressed by:
${HR}_{x}\ =N(\mu,\sigma^{2})$ (9)
Where $N$ represents a normal distribution, and $x$ denotes the ground truth or the prediction. And $\mu$ denotes the mean of the distribution, where $\mu\ =\ maxIdx(PSD({BVP}_{x}\ ))$. The $maxIdx$ indicates the index of the maximum value, and $PSD$ indicates the power spectrum. The HR distance is expressed by:
$\mathcal{L}_{HR}\ =KL({HR}_{gt},{HR}_{pred})$ (10)
@@ 3.4 HR Hybrid Loss
The temporal loss $\mathcal{L}_{Time}$ is computed using the negative Pearson correlation coefficient, and the frequency loss is calculated using the cross-entropy between the predicted BVP wave’s frequency domain and the heart rate derived from the actual BVP wave. The frequency loss $\mathcal{L}_{Freq}$ is then expressed by:
$\small\mathcal{L}_{Freq}=CE(maxIdx(PSD({BVP}_{gt})),PSD({BVP}_{pred}))$ (11)
The overall loss is expressed by:
$\mathcal{L}_{overall}=\alpha\ast\mathcal{L}_{Time}+\beta\ast\ \mathcal{L}_{Freq}+\gamma\ast\mathcal{L}_{HR}$ (12)
Where $\alpha=0.2$, $\beta=1.0$, $\gamma=1.0$ in our experiment.
@@ 4.1 Dataset and performance metric
Three publicly available datasets were used for evaluation: PURE [40], UBFC-rPPG [2], and MMPD [41]. Specifically, PURE contains 60 1-minute videos comprised of 10 participants engaged in six activities, with a video resolution of 640x480 and a frame rate of 30Hz. UBFC-rPPG contains 42 videos, recording subjects in a stationary state, with a resolution of 640x480 and a frame rate of 30Hz. MMPD comprised 33 participants with Fitzpatrick skin types 3-6. These participants performed four different activities (static, head rotation, speech, and walking) under four lighting conditions (LED-high, LED-low, incandescent, natural). The dataset included 660 1-minute videos with a resolution of 320x240 and a frame rate of 30Hz. Notably, MMPD has two versions: a compressed mini-MMPD and an uncompressed MMPD. In this study, the mini-MMPD was chosen for experimentation. Clearly, among the three datasets, PURE and UBFC are relatively easy, while MMPD stands out with a significantly larger sample size and more complex environment. Five commonly used metrics: Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), Mean Absolute Percentage Error (MAPE), Pearson correlation coefficient ($\rho$), and Signal-to-Noise Ratio (SNR) were used for evaluation.
@@ 4.2 Implementation Details
We leveraged the rPPG toolbox [24], an open-source toolbox based on PyTorch, to implement the proposed method. A comparative analysis was conducted against several state-of-the-art methods, all implemented with the same toolbox. We followed the protocol outlined in [53] by employing random upsampling, downsampling, and horizontal flipping for data augmentation. For each video segment, the facial region was cropped and resized to 128×128 pixels in the first frame, and this region was fixed in the subsequent frames. In post-processing, a second-order Butterworth filter was applied to filter the predicted PPG waveform. The Welch algorithm was then employed to compute the power spectral density for further heart rate estimation. One more thing to note is that due to suboptimal performance on most datasets with PhysFormer’s original configuration, similar to the treatment of PhysNet[51] in the toolbox, we utilized the normalized frame difference as the input for PhysFormer. For a fair comparison, all experiments were conducted under the same experimental setting. Default hyperparameters from the toolbox were employed. The learning rate was set to 9e-3. The batch size was set to 4, and the training process spanned 30 epochs. Network training was conducted on an NVIDIA RTX A6000.
Table 1: Intra-dataset results on PURE and UBFC.
PURE UBFC
Method MAE RMSE $\rho$ MAE RMSE $\rho$
HR-CNN[39] 1.84 2.37 0.98 4.90 5.89 0.64
DeepPhys[3] 0.83 1.54 0.99 6.27 10.82 0.65
SynRhythm[30] 2.71 4.86 0.98 5.59 6.82 0.72
PhysNet[51] 2.10 2.60 0.99 2.95 3.67 0.97
rPPGNet[52] 0.74 1.21 - 0.56 0.73 0.99
Meta-rPPG[17] 2.52 4.63 0.98 5.97 7.42 0.53
TS-CAN[22] 2.48 9.01 0.92 1.70 2.72 0.99
Siamese[42] 0.51 1.56 0.83 0.48 0.97 -
PulseGAN[38] 2.28 4.29 0.99 - - -
Dual-GAN[27] 0.82 1.31 0.99 0.44 0.67 0.99
AND-rPPG[26] - - - 2.67 4.07 0.92
TDM[6] 1.83 2.30 0.99 2.32 3.08 0.99
PhysFormer[56] 1.10 1.75 0.99 0.50 0.71 0.99
EfficientPhys[23] - - - 1.14 1.81 0.99
RADIANT[13] - - - 2.91 4.52 -
Li et al.[20] 1.44 2.50 - 0.76 1.62 -
LSTC[18] - - - 0.70 1.00 0.99
zhang et al.[59] 1.53 2.29 0.99 - - -
Ours 0.27 0.47 0.99 0.50 0.78 0.99
Table 2: Intra-dataset results on MMPD.
MMPD
Method MAE RMSE MAPE $\rho$ SNR
DeepPhys[3] 22.27 28.92 23.9 -0.03 -13.44
PhysFormer[56] 11.99 18.41 12.50 0.18 -9.51
EfficientPhys[23] 13.47 21.32 14.19 0.21 -9.25
TS-CAN[22] 9.71 17.22 10.36 0.44 -6.69
PhysNet[51] 4.80 11.80 4.74 0.60 1.77
Ours 3.07 6.81 3.24 0.86 5.46
@@ 4.3 Intra-dataset evaluation
For the evaluation on the PURE dataset, we adhered to the protocol outlined in [39, 27]. Specifically, the first 60% of the dataset was utilized for training, while the remaining 40% was allocated for testing. Similarly, for the evaluation on the UBFC dataset, we followed the protocol specified in [38, 27], selecting the initial 30 samples for training and the subsequent 12 samples for testing. As illustrated in Table 1, on the PURE dataset, our method outperformed state-of-the-art methods across all metrics. In the case of the UBFC dataset, our method demonstrated comparable performance to state-of-the-art methods.
Since the relative simplicity of PURE and UBFC, the performance of state-of-the-art methods on these two datasets is about to be saturated. We further conducted an evaluation with the challenging MMPD dataset, partitioned into training, validation, and test sets with a ratio of 7:1:2. As depicted in Table 2, we compared our method against existing state-of-the-art approaches. Note that the MMPD dataset is newly published. The performance of existing methods on this dataset is reimplemented by ourselves with the rPPG toolbox. Confronting the more complex environmental conditions of MMPD, our approach achieved the lowest MAE (3.07), RMSE (6.81), and the highest $\rho$ (0.86), surpassing previous approaches by a significant margin. These showcased superior rPPG feature extraction capabilities, demonstrating the enhanced robustness of RhythmFormer.
The observed phenomenon of other Transformer-based methods performing well on PURE and UBFC but showing a large performance margin on MMPD suggests that: in more complex environments, coarse-grained extraction of rPPG features is highly susceptible to noise interference, leading to poor robustness. RhythmFormer leveraging periodic sparse attention for fine-grained extraction of rPPG features, demonstrates its potential for applications in complex real-world environments with outstanding robustness.
Table 3: Cross-dataset results.

### v3

@@ 3.3 Hierarchical Temporal Periodic Transformer
Previous studies [34, 35] have demonstrated the effectiveness of modeling periodic tasks with multiple temporal scales, primarily extracting temporal information through concatenation from paths at various scales. In the rPPG field, similar to many spatial multi-scale methods [26, 27], we hypothesize that a pyramid-shaped hierarchical multi-scale learning may facilitate the provision of information from earlier learned scales to later learned scales.
The proposed hierarchical temporal periodic transformer is composed of three TPT blocks with downsampling factors of $n$, where the values of $n$ are $1,2,3$ respectively. It first reinforces inter-frame rPPG features at a smaller temporal scale and then models periodic features at a larger temporal scale, effectively reducing noise interference. Specifically, for the TPT block with a downsampling factor of n, the temporal downsampling module aggregates the temporal information at a granularity of $2^{n}$ without overlap. The downsampling module consists of a BN layer and a 3D convolution layer with the kernel and stride size of $(2\times 1\times 1)$, resulting in an output $X_{sample}\in\mathbb{R}^{C\times{T}/{2^{n}}\times H/16\times W/16}$. The first TPT block does not have a dedicated downsampling module; instead, the preceding patch embedding performs the function of the downsampling module.
@@
The first part of the TPT block is the periodic sparse attention module. Figure 2 illustrates the attention mechanism, where a pre-attention stage is added before the attention to ensure sufficient computational resources for fine-grained operations. The specific computation steps are shown in the blue part of Figure 3. First, the input to the module is projected. Here, Q and K are processed using the temporal difference convolution (TDC) module, while a linear projection is used to project V. The TDC module [36] has been proven to capture fine-grained local temporal difference features effectively [37, 20, 11, 32]. Then the projected Q and K are subjected to a 3D average pooling operation to compute attention within a larger spatiotemporal receptive field, which filters out low-correlation key-value pairs. Specifically, for input $Q,K\in\mathbb{R}^{C\times{T}/{2^{n}}\times H/16\times W/16}$, the pooling convolutional kernel size is $(T^{\prime}\times H^{\prime}\times W^{\prime})$.
$T^{\prime}=[\frac{T}{T_{s}}],H^{\prime}=[\frac{H/16}{4}],W^{\prime}=[\frac{W/16}{4}].$ (3)
@@
Where $T_{s}=MAX(2^{x},2^{n})$, and $x$ is the partition coefficient. In the experiments, since the frame rate of the dataset is uniformly downsampled to 30, $x$ is set to 2. The output $Q^{\prime}$ and $K^{\prime}$ after pooling are used to compute the attention in the pre-attention stage. The scores of the pre-attention are computed as follows:
$Score\ =\ Q^{\prime}(K^{\prime})^{\top}.$ (4)
@@
The pre-attention stage is designed to obtain the indices of top-k tokens with high correlation, to filter out low-correlation key-value pairs. The pre-attention captures spatiotemporal correlations within a large receptive field, gathering the regions highly correlated with Q in both K and V. If a token from the facial area is given as a query, it corresponds to gathering facial regions with similar temporal phases to the query. The gathered key and value are expressed as:
$K^{g}\ =\ gather(K,topkIndex(Score)),$ (5)
$V^{g}\ =\ gather(V,topkIndex(Score)).$
@@
The refined attention stage is computed on the top-k (high-correlation) regions identified after the pre-attention gathering. Subsequently, the local positional information is enhanced through a local context enhancement (LCE) module [29, 28]. LCE is a 3D convolution layer with a kernel size of $(3\times 3\times 3)$. The $i_{th}$ self-attention head output ${SA}_{i}$ is:
${SA}_{i}\ =Softmax(\ Q{(K^{g})}^{\top}/\sqrt{d})\ V^{g}\ +\ LCE(V).$ (6)
@@
Where $d$ represents the dimensionality of single-head attention. The output of multi-head self-attention (MHSA) is obtained by concatenating the outputs from all h attention heads, followed by a linear projection through the weight matrix $W\in\mathbb{R}^{d\times d}$.
$MHSA=Concat(SA_{1},SA_{2},...,SA_{h})W.$ (7)
@@
Like vanilla transformer [25], residual connection, BN, and feed-forward would be conducted after MHSA. Diverging from that, inspired by [11, 32], an additional hidden layer is incorporated into the feed-forward layer. This hidden layer consists of a 3D convolution layer, BN, and GELU, intended to refine local inconsistency and reinforce relative position cues.
3.4 Model Training
We employed a hybrid loss function incorporating both time-domain and frequency-domain components [37], which can be expressed as $\mathcal{L}_{overall}=0.2\cdot\mathcal{L}_{Time}+\mathcal{L}_{Freq}$. Specifically, the temporal loss $\mathcal{L}_{Time}$ is computed using the negative Pearson correlation coefficient, and can be represented as:
$\mathcal{L}_{Time}=1-\frac{\sum_{i=1}^{n}\left({S}_{gt_{i}}-\overline{{S}_{gt}}\right)\left({S}_{pred_{i}}-\overline{{S}_{pred}}\right)}{\sqrt{\sum_{i=1}^{n}\left({S}_{gt_{i}}-\overline{{S}_{gt}}\right)^{2}}\sqrt{\sum_{i=1}^{n}\left({S}_{pred_{i}}-\overline{{S}_{pred}}\right)^{2}}}.$ (8)
@@ 3.4 Model Training
Where $S$ denotes the Blood Volume Pulse waves. The frequency loss is calculated using the cross-entropy between the predicted BVP wave’s frequency domain and the heart rate derived from the ground truth BVP wave. The frequency loss $\mathcal{L}_{Freq}$ is then expressed by:
$\mathcal{L}_{Freq}=CE(maxIndex(PSD({S}_{gt})),PSD({S}_{pred})).$ (9)
Where $CE$ denotes cross-entropy, $maxIndex$ indicates the index of the maximum value, and $PSD$ indicates the power spectrum.
On this basis, we further employed an HR constraint, denoted as $B_{HR}$ which serves as a bias term for selecting the optimal checkpoint. It is calculated using the distance between the ground truth HR distribution and the predicted HR distribution. The HR distribution can be represented as $HR=N(\mu,\sigma^{2})$. Where $N$ represents the normal distribution, and $\mu=maxIndex(PSD(S))$. The HR constraint is expressed by $B_{HR}\ =KL({HR}_{gt},{HR}_{pred})$. $KL$ denotes the Kullback-Leibler divergence. The final checkpoint selection is performed by $\mathcal{L}_{overall}+B_{HR}$.
@@ 4.1 Dataset and Performance Metric
Five publicly available datasets were used for evaluation: PURE [38], UBFC-rPPG [39], COHFACE [40], VIPL-HR [16], MMPD [41]. Specifically, PURE contains 60 1-minute videos comprised of 10 participants engaged in six activities, with a video resolution of 640$\times$480 and a frame rate of 30 fps. UBFC-rPPG contains 42 videos, recording subjects in a stationary state, with a resolution of 640$\times$480 and a frame rate of 30 fps. COHFACE consists of 160 one-minute video sequences from 40 participants (12 females and 28 males), with a frame rate of 20 fps. VIPL-HR includes 2,378 RGB videos from 107 participants, captured using three RGB cameras, with an unstable fps. MMPD comprised 33 participants with Fitzpatrick skin types 3-6. These participants performed four different activities (static, head rotation, speech, and walking) under four lighting conditions (LED-high, LED-low, incandescent, natural). The dataset included 660 1-minute videos with a resolution of 320$\times$240 and a frame rate of 30 fps. Notably, MMPD has two versions: a compressed mini-MMPD and an uncompressed MMPD. In this study, the mini-MMPD was chosen for experimentation. Among the five datasets, PURE and UBFC-rPPG are relatively easy, while MMPD and VIPL-HR stand out with a significantly larger sample size and more complex environment. Five commonly used metrics: Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), Mean Absolute Percentage Error (MAPE), Pearson correlation coefficient ($\rho$), and Signal-to-Noise Ratio (SNR) were used for evaluation.
@@ 4.2 Implementation Details
We leveraged the rPPG toolbox [42], an open-source toolbox based on PyTorch, to implement the proposed method. A comparative analysis was conducted against several state-of-the-art methods, all implemented with the same toolbox. For each video segment, the facial region was cropped and resized to 128$\times$128 pixels in the first frame, and this region was fixed in the subsequent frames. In post-processing, a second-order Butterworth filter (cutoff frequencies: 0.75Hz and 2.5 Hz) was applied to filter the predicted PPG waveform. The Welch algorithm was then employed to compute the power spectral density for further heart rate estimation. The value of k for top-k was set to 40. The batch size was set to 4, and the training process spanned 30 epochs. The experiments were performed using a computer that includes an AMD EPYC 7542 32-Core Processor. Network training was conducted on NVIDIA RTX A6000.
Figure 5: Heart rate distribution among different datasets.
4.3 Details of Data Augmentation
As shown in figure 5, the HR distribution among different datasets is highly imbalanced. Hence, we employed a random sampling data augmentation strategy [37]. For cases where the heart rate, calculated from the ground truth BVP wave, exceeded 90 beats per minute (bpm), random downsampling was applied; for heart rates below 75 bpm, random upsampling was performed. Additionally, random horizontal flips were applied to all videos.
@@ 4.4 Intra-dataset Evaluation
Estimation on PURE and UBFC. For the evaluation of the PURE dataset, we adhered to the protocol outlined in [15, 30]. Specifically, the first 60% of the dataset was utilized for training, while the remaining 40% was allocated for testing. For the evaluation of the UBFC-rPPG dataset, we followed the protocol specified in [43, 30], selecting the initial 30 samples for training and the subsequent 12 samples for testing. According to these protocols, none of these datasets have a validation set, and the last epoch of training is used for testing. As illustrated in Table 1, on the PURE dataset, our method outperformed state-of-the-art methods across all metrics. In the case of the UBFC-rPPG dataset, our method demonstrated comparable performance to state-of-the-art methods.
Estimation on COHFACE. We adhered to the protocol outlined in [40]. Specifically, the first 60% of the dataset was utilized for training, while the remaining 40% was allocated for testing. As illustrated in Table 1, our method outperforms previous approaches, achieving an MAE of 1.17. For the heavily compressed COHFACE dataset, the rPPG features may be weaker, so we conducted further exploration. Table 2 presents the results using raw frames as input, processed through different stem modules. As seen in the table, the fusion stem effectively enhances rPPG features and guides attention, leading to significantly better performance with both PhysFormer and RhythmFormer.
PURE UBFC COHFACE VIPL-HR MMPD
[…]
HR-CNN [15] 1.84 2.37 0.98 4.90 5.89 0.64 8.10 10.78 0.29 - - - - - -
DeepPhys [2] 0.83 1.54 0.99 6.27 10.82 0.65 6.89 13.89 0.34 11.00 13.80 0.11 22.27 28.92 -0.03
SynRhythm [44] 2.71 4.86 0.98 5.59 6.82 0.72 - - - - - - - - -
RhythmNet [16] - - - - - - - - - 5.30 8.14 0.76 - - -
PhysNet [19] 2.10 2.60 0.99 2.95 3.67 0.97 5.38 10.76 - 10.80 14.80 0.20 4.80 11.80 0.60
Meta-rPPG [45] 2.52 4.63 0.98 5.97 7.42 0.53 9.31 12.27 0.19 - - - - - -
TS-CAN [17] 2.48 9.01 0.92 1.70 2.72 0.99 - - - - - - 9.71 17.22 0.44
CVD [46] 1.29 2.01 0.98 2.19 3.12 0.99 14.50 16.50 0.30 5.02 7.97 0.79 - - -
DeeprPPG [47] 0.28 0.43 0.99 0.67 1.70 0.99 3.07 7.06 0.86 - - - - - -
PulseGAN [43] 2.28 4.29 0.99 1.19 2.10 0.98 - - - - - -
Dual-GAN [30] 0.82 1.31 0.99 0.44 0.67 0.99 - - - 4.93 7.68 0.81 - - -
AND-rPPG [48] - - - 2.67 4.07 0.92 6.81 8.06 0.63 - - - - - -
ETA-rPPGNet [49] 0.34 0.77 0.99 1.46 3.97 0.93 4.67 6.65 0.77 - - - - - -
PhysFormer [11] 1.10 1.75 0.99 0.50 0.71 0.99 16.17 17.92 0.19 4.97 7.79 0.78 11.99 18.41 0.18
EfficientPhys [23] - - - 1.14 1.81 0.99 - - - - - - 13.47 21.32 0.21
TFA-PFE [50] 1.44 2.50 - 0.76 1.62 - 1.31 3.92 - - - - - - -
SINC [51] 0.61 1.84 0.99 0.59 1.83 0.99 - - - - - - - - -
Li et al. [52] 0.64 1.16 0.99 0.48 0.64 0.99 - - - 5.19 8.26 0.78 - - -
Yue et al. [22] 1.23 2.01 0.99 0.58 0.94 0.99 - - - - - - - - -
PhysFormer++ [32] - - - - - - - - - 4.88 7.62 0.80 - - -
Face2PPGbest [9] 1.20 1.80 - 0.90 2.80 - 7.50 9.80 - - - - - - -
Contrast-Phys+ [21] 0.48 0.98 0.99 0.21 0.80 0.99 - - - - - - - - -
Ours 0.27 0.47 0.99 0.50 0.78 0.99 1.17 3.36 0.97 4.51 7.98 0.78 3.07 6.81 0.86
Table 1: Intra-dataset results on PURE, UBFC-rPPG, COHFACE, VIPL-HR, and MMPD. The best results are in bold.
PhysFormer RhythmFormer
[…]
Vanilla Stem 16.17 17.92 25.93 0.19 -11.42 2.05 5.36 3.33 0.91 4.76
Fusion Stem 2.55 5.81 4.18 0.90 -5.28 1.17 3.36 1.84 0.97 9.28
Table 2: Intra-dataset results trained on COHFACE.
Estimation on VIPL-HR and MMPD. The performance of state-of-the-art methods on the previous three datasets is nearing saturation. We further conducted an evaluation using the challenging dataset. For the VIPL-HR dataset, we followed the subject-exclusive 5-fold cross-validation protocol, as outlined in [16, 11]. For the MMPD dataset, it is partitioned in sequence into training, validation, and test sets with a ratio of 7:1:2. As depicted in Table 1, we compared our method against existing state-of-the-art approaches. Note that the MMPD dataset is newly published. The performance of existing methods on this dataset is reimplemented by ourselves with the rPPG toolbox. Confronted with more complex environmental conditions, our method achieves optimal performance on the MMPD dataset and shows comparable performance to state-of-the-art methods on the VIPL-HR dataset. The reason for the lower MAE but higher RMSE on VIPL-HR may be due to the presence of some extremely challenging examples in the dataset, where the head movement speed is very fast and the average frame rate is below 20. In such cases, fine-grained extraction may not provide an advantage.
To further examine the correlation between the predicted HR and the ground truth HR, we provide scatter plots and Bland-Altman plots of intra-dataset evaluation on MMPD in Figure 6 and Figure 7. As depicted in these figures, compared to other methods, the HR estimated by RhythmFormer shows a strong correlation with the Ground Truth HR. The scatter plot is generally aligned with the y = x line, and the points in the Bland-Altman plot mostly fall within the confidence interval, indicating the effectiveness and robustness of RhythmFormer. Meanwhile, we have noticed the presence of outliers, primarily distributed in the heart rate range of 100-110 bpm. This may be due to the distribution differences in the MMPD dataset, where the data count of heart rates exceeding 100 bpm sharply decreases (see Figure 5).
Additionally, examples of the time-domain and frequency-domain plots of pulse waves predicted by RhythmFormer are provided, as illustrated in Figure 8. The comparison of the raw rPPG waves, predicted rPPG signals and ground truth PPG signals in the time domain provides a direct evaluation of the model’s ability to remove noise and artifacts, thereby intuitively demonstrating the model’s effectiveness. The predicted signals exhibit a strong alignment with the ground truth signals in terms of trends, yet they lack detailed high-frequency information. Notably, our model robustly captures the peaks and variations of BVP signals, serving as the core judgment basis for predicting heart rate from video data.
Figure 6: Scatter plots from intra-dataset evaluation on MMPD using DeepPhys (a), PhysNet (b), TS-CAN (c), PhysFormer (d), EfficientPhys (e) and RhythmFormer (f).
Figure 7: Bland-Altman plots from intra-dataset evaluation on MMPD using DeepPhys (a), PhysNet (b), TS-CAN (c), PhysFormer (d), EfficientPhys (e) and RhythmFormer (f).
Figure 8: Examples of HR estimation on PURE (a), UBFC-rPPG (b) and MMPD (c) using RhythmFormer. “Raw” represents the raw rPPG waves (the spatial average of video frames).
4.5 Cross-dataset Evaluation
We followed the protocol outlined in [42] for cross-dataset evaluation. The models were trained on PURE and UBFC-rPPG individually and subsequently tested on PURE, UBFC-rPPG, and MMPD datasets. The training dataset was split, with the initial 80% used for training and the remaining 20% for validation. Comparative results against state-of-the-art methods are presented in Table 3. The results highlight RhythmFormer’s superior capability in modeling domain-invariant rPPG features and its generalization across diverse datasets. This suggests that fine-grained feature extraction may more effectively mitigate interference from background noise and personalized facial features, enabling the model to learn domain-invariant features.
Figure 9: Scatter plots and Bland-Altman plots of results in cross-dataset evaluation. (a, b) is training on PURE and testing on UBFC-rPPG; (c, d) is training on UBFC-rPPG and testing on PURE.